Exercise 1: Consider an Australian financial institution with a trading position of 1’000’000 EUR. The current exchange rate is AUD 1.6 / EUR.

a) Calculate the value of the position in AUD.
b) The volatility of the exchange rate is AUD 0.1 / EUR. Calculate the volatility in percent of the exchange rate.

c) Calculate the extreme adverse change in the foreign exchange rate for α = 1% in percent and in AUD x / EUR.

d) Calculate the daily earnings at risk for the foreign exchange rate exposure. e) Calculate the 5-day earnings at risk for the foreign exchange rate exposure.
Answer #1

a. Value of position in AUD = Value in EUR * Exchange rate per EUR = 1,000,000 * 1.6 = 1,600,000

b. Volatility in percentage = (0.1 *100) / 1.6 = 6.25%

c. Extreme change in rate if alpha = 1% , is 1% and exchange rate 1% * 1.6 = 0.016

d & e.. Need more information like movement in exchange rate per to calculate